OpenStudy (anonymous):

Which of the following is the solution to the equation 8^(y - 2) = 2y

OpenStudy (anonymous):

y = 1 y = 3 y = -3 y = -1

sam (.sam.):

\[\Huge 8^{y - 2} = 2^y~~~~~,~~~~~\text{Use}~~8=2^2\] \[\Huge 2^{2(y - 2)} = 2^y\] Since the base are same, you can bring down the exponents 2(y-2)=y Try to solve from here?

OpenStudy (anonymous):

8(y-2)= 2y 8y-16=2y 6y=16 y =8/3 ??

OpenStudy (anonymous):

i get an answer but its not one of the choices

sam (.sam.):

Did you follow the steps?

OpenStudy (anonymous):

I always come out with y=8/3 64/3 - 48/3 =16/3 16/3 = 16/3 it checks

OpenStudy (anonymous):

or im confused haha

sam (.sam.):

You need to change the 8 to 2, \[\Huge 8^{y - 2} = 2^y~~~~~,~~~~~\text{Use}~~8=2^3\] \[\Huge 2^{3(y - 2)} = 2^y\] Then 3(y-2)=y

sam (.sam.):

3y-6=y 2y=6 y=3

OpenStudy (anonymous):

thank you